# Parameters Optimization in Trajectory Planning Using Diffrentiable Convex Programing

### v1

@@ 3 Differentiable Sequencial Convex Programing
3.1 Differentiable Second-Order Cone Programing
@@
Constraints with indices $j\notin\mathcal{A}_{\mathrm{lin}}$ and $i\notin\mathcal{A}_{\mathrm{soc}}$ do not contribute to the stationarity condition or subsequent derivations. In the remainder of this section, all summations are taken only over active constraints. Based on complementary slackness, one obtains
@@
where the last inequality follows from the Cauchy–Schwarz inequality. In addition, since both $\boldsymbol{\lambda}_{i}$ and $\boldsymbol{g}_{i}$ belong to the cone $\mathcal{Q}_{m_{i}}$, we have
@@
Combining Eq.(27) through Eq.(32), we obtain the linear system
@@ 3.2.1 Forward process of differentiable sequential convex programming
Using a numerical interior-point solver [12], we obtain the optimal solution of the subproblem, denoted $\boldsymbol{z}^{(k)}_{\star}$. This solution serves simultaneously as the reference trajectory for the next iteration $\boldsymbol{z}_{\mathrm{ref}}^{(k+1)}=\boldsymbol{z}^{(k)}_{\star}$. The SCP forward process is terminated once the improvement in the objective falls below the prescribed threshold given by
@@ 3.2.2 Backward process of differentiable sequential convex programming
$\mathcal{L}=\mathcal{L}\left(\boldsymbol{z}_{\mathrm{ref}}^{(0)},\boldsymbol{z}_{\mathrm{ref}}^{(1)},\dots,\boldsymbol{z}_{\mathrm{ref}}^{(K)},\boldsymbol{\theta}_{\mathrm{SCP}},\boldsymbol{\theta}_{\mathrm{TP}}\right)=\mathcal{L}\left(\boldsymbol{\theta}^{(1)},\dots,\boldsymbol{\theta}^{(K)},\boldsymbol{\theta}^{(K+1)}\right)$ (39)
where $\boldsymbol{\theta}^{(K+1)}=\left(\boldsymbol{z}_{\star}^{(K)},\boldsymbol{\theta}_{\mathrm{SCP}},\boldsymbol{\theta}_{\mathrm{TP}}\right)$. The cost function thus depends explicitly on the optimal solution of the final iteration as well as implicitly on all intermediate iterations through the sequence of reference trajectories. Since the SCP process is parameterized by $\boldsymbol{\theta}_{\mathrm{SCP}},\boldsymbol{\theta}_{\mathrm{TP}}$, the chain rule gives
$\frac{\partial\mathcal{L}}{\partial\boldsymbol{\theta}_{\mathrm{SCP}}}$ $=\sum_{k=1}^{K+1}\frac{\partial\mathcal{L}}{\partial\boldsymbol{\theta}^{(k)}}\frac{\partial\boldsymbol{\theta}^{(k)}}{\partial\boldsymbol{\theta}_{\mathrm{SCP}}}$ (40)
$\frac{\partial\mathcal{L}}{\partial\boldsymbol{\theta}_{\mathrm{TP}}}$ $=\sum_{k=1}^{K+1}\frac{\partial\mathcal{L}}{\partial\boldsymbol{\theta}^{(k)}}\frac{\partial\boldsymbol{\theta}^{(k)}}{\partial\boldsymbol{\theta}_{\mathrm{TP}}}$
The derivative of the cost function with respect to the final optimal solution $\boldsymbol{z}_{\star}^{(K)}$ can be computed directly from the definition of the loss, yielding $\frac{\partial\mathcal{L}}{\partial\boldsymbol{\theta}^{(K+1)}}$. For an intermediate iteration $k$, the gradient with respect to the parameter set $\boldsymbol{\theta}^{(\theta)}$ is given by
$\frac{\partial\mathcal{L}}{\partial\boldsymbol{\theta}^{(k)}}=\frac{\partial\mathcal{L}}{\partial\boldsymbol{z}_{\star}^{(k)}}\frac{\partial\boldsymbol{z}_{\star}^{(k)}}{\partial\boldsymbol{\theta}^{(k)}},k\in[1,K]$ (41)
The Jacobian $\frac{\partial\boldsymbol{z}_{\star}^{(k)}}{\partial\boldsymbol{\theta}^{(k)}}$ is obtained from the linear sensitivity equation Eq.(33). However, computing this Jacobian explicitly would require forming blockwise inverses of $\boldsymbol{H}^{(k)}$ and evaluating explicit analytic derivatives of matrices such as $\boldsymbol{A}^{(k)}$ with respect to $\boldsymbol{\theta}^{(k)}$. To avoid this computational burden, an auxiliary linear system is introduced as
@@
Here, $\boldsymbol{P}^{(k)}$ is a permutation matrix obtained using an approximate minimum-degree method, $\boldsymbol{L}^{(k)}$ is lower triangular, and $\boldsymbol{D}^{(k)}$ is a diagonal matrix. A small regularization term is added to the diagonal of $\boldsymbol{H}^{(k)}$ to ensure numerical stability. Since the coefficient matrices arising in SCP are typically large-scale and sparse, exploiting sparse symmetric matrix factorizations to solve the associated linear systems is computationally advantageous and significantly accelerates CPU-based computations. Taken together, Eq.(33) and Eq.(42) form a coupled forward-backward sensitivity propagation system. Eliminating intermediate variables leads to the differential of the loss
@@
$\frac{\partial\mathcal{L}}{\partial\boldsymbol{b}}=\boldsymbol{y}_{\boldsymbol{\nu}},\,\,\frac{\partial\mathcal{L}}{\partial\boldsymbol{h}_{0+}}=\boldsymbol{y}_{\boldsymbol{\mu}},\,\,\frac{\partial\mathcal{L}}{\partial b_{i}}=y_{s,i},\,\,\frac{\partial\mathcal{L}}{\partial\boldsymbol{h}_{i}}=-\lambda_{i,0}\left(\boldsymbol{I}-\frac{\boldsymbol{g}_{i,1}\boldsymbol{g}_{i,1}^{\top}}{g_{i,0}^{2}}\right)\frac{\boldsymbol{G}_{i}\boldsymbol{y}_{\boldsymbol{z}}}{g_{i,0}}-y_{s,i}\frac{\boldsymbol{g}_{i,1}}{g_{i,0}}$
@@
Eq.(40), Eq.(45), and Eq.(46) describe the complete backpropagation pathway of the differentiable SCP framework.
@@ 4 Simulations and Results
To demonstrate the generality and practical utility of the proposed DSCP framework, three numerical experiments are conducted across two representative trajectory planning scenarios: powered powered guidance (PDG) for reusable rockets and entry flight for hypersonic gliding vehicles (HGVs). These experiments target three different categories of parameters, including nonconvex decision variables, hyperparameters of the SCP procedure, and vehicle-design parameters. All simulations are executed in Python 3.8.20. Gradient backpropagation is implemented using PyTorch 1.12.0, and the convex subproblems are solved using the ECOS solver[12]. As ECOS currently supports only CPU computation, all simulations are performed on the CPU.
@@ 4.1 Optimization of Terminal Time in Powered Descent Guidance
This subsection applies the DSCP framework to a classical powered descent guidance (PDG) problem[7]. The objective is to learn and optimize the terminal time, which is a nonconvex and problem-dependent variable that influences fuel optimality.
@@ 4.1.1 Powered Descent Guidance for Reusable Rockets
$\dot{\boldsymbol{v}}=\frac{\boldsymbol{T}}{m}+\boldsymbol{g}_{E}$
$\dot{m}=-\frac{\left\|\boldsymbol{T}\right\|_{2}}{I_{\mathrm{sp}}g_{0}}$
The state vector is $\boldsymbol{x}_{\mathrm{PDG}}=\left[\boldsymbol{r}^{\top},\boldsymbol{v}^{\top},m\right]^{\top}$, where $\boldsymbol{r}$ and $\boldsymbol{v}$ denote the position and velocity vectors, and $m$ is the rocket mass. The parameter $I_{\mathrm{sp}}$ denotes the specific impulse, and $\boldsymbol{g}_{E}=[g_{0},0,0]^{\top}$ is the gravitational acceleration vector, where $g_{0}$ is the standard gravitational constant. The descent guidance is subject to the following operational constraints:
$\boldsymbol{x}(t_{0})=\boldsymbol{x}_{0}$ (48)
@@
$T_{\min}\leq\left\|\boldsymbol{T}\right\|_{2}\leq T_{\max}$
@@
$\sqrt{T_{y}^{2}+T_{z}^{2}}-\tan{\eta_{\max}}T_{x}\leq 0$
where $T_{\mathrm{min}}$ and $T_{\mathrm{max}}$ are thrust bounds, $\beta_{\max}$ is the maximum allowable glideslope angle, and $\eta_{\max}$ is the maximum thrust pointing angle. The performance index for fuel-optimal descent is
@@
Fuel-optimal powered descent is inherently a free-final-time problem. The present experiment evaluates the DSCP framework’s ability to learn and optimize such nonconvex timing parameters. The work in [7] converts the free-final-time problem into multiple fixed-time problems and performs a grid search over the terminal time, followed by training a neural network to approximate the optimal relationship. In contrast, DSCP uses the gradients obtained by backpropagation through the SCP iterations, allowing direct end-to-end training of a neural network without terminal-time enumeration.
@@ 4.1.2 Verification of Differentiability
This experiment evaluates the accuracy of the gradients produced by DSCP. A virtual-control-augmented SCP method is used to solve the fixed-time fuel-optimal PDG problem. Let $K_{\mathrm{PDG}}$ be the total number of SCP iterations and $N_{\mathrm{PDG}}$ be the number of discretization intervals. The scalar loss used for backpropagation is
@@
Simulation results are shown in Fig. 2 through Fig. 3. Figure 2(a) illustrates the relationship between terminal time and maximum remaining mass for a fixed initial condition. The results indicate that no feasible solution exists within the interval $32.00,32.05]\mathrm{s}$, and solutions in $32.05,32.10]\mathrm{s}$ exhibit abnormal behavior, where virtual control terms do not converge to zero. This failure occurs because the terminal time is too short for feasible descent. When $t_{\mathrm{f}}>32.10\mathrm{s}$, the SCP iterations converge reliably. With two-decimal precision, the optimal terminal time is $32.81\mathrm{s}$, consistent with the result reported in [7].
@@ 4.1.3 The Training of Neural Network for Predicting the Terminal Time
$\boldsymbol{r}(t_{0})=[5200\pm 200,\pm 600,\pm 600]^{\top}(\mathrm{m}),\boldsymbol{v}(t_{0})=[-150\pm 15,\pm 30,\pm 30]^{\top}(\mathrm{m/s})$ (52)
@@
Figure 4 displays the training behavior, including the learning-rate profile and the average absolute value of the output gradient per epoch. As training progresses, gradient magnitudes decrease steadily, indicating improved numerical stability. Test results are presented in Figs.5 and 6. Under a representative -vertical velocity-altitude profile with $r_{y}=r_{z}=100\mathrm{m},v_{y}=v_{z}=0$, the maximum relative prediction error is below $3\times 10^{-4}$. For a horizontal profile with $r_{x}=5200\mathrm{m},v_{x}=-135\mathrm{m/s},v_{y}=v_{z}=0$, the maximum relative error remains below $4\times 10^{-4}$. These results confirm that the proposed DSCP framework can be seamlessly integrated with neural networks and can effectively train them to learn nonconvex variables such as the optimal terminal time in trajectory planning problems.
@@ 4.2 Trust-Region Penalty Coefficient Optimization
The trust-region penalty coefficients constitute a set of critical design parameters in construction of the convexified subproblems of SCP. This subsection considers the entry trajectory planning of hypersonic gliding vehicles (HGVs) and investigates the effects of the magnitudes of the trust-region penalty coefficients on the convergence rate and numerical behavior of the trajectory optimization process. In contrast to adaptive trust-region updating strategies used in previous studies, the present investigation focuses on fixed penalty values that remain constant throughout all SCP iterations.

### v2

@@ 3 Differentiable Sequential Convex Programming
3.1 Differentiable Second-Order Cone Programming
@@
Constraints with indices $j\notin\mathcal{A}_{\mathrm{lin}}$ and $i\notin\mathcal{A}_{\mathrm{soc}}$ do not contribute to the stationarity condition or subsequent derivations [panBPQPDifferentiableConvex]. In the remainder of this section, all summations are taken only over active constraints. Based on complementary slackness, one obtains
@@
where the last inequality follows from the Cauchy-Schwarz inequality. In addition, since both $\boldsymbol{\lambda}_{i}$ and $\boldsymbol{g}_{i}$ belong to the cone $\mathcal{Q}_{m_{i}}$, it holds that
@@
Combining Eq.(27) through Eq.(32), it yields the linear system
@@ 3.2.1 Forward Process of Differentiable Sequential Convex Programming
Using a numerical interior-point solver [domahidiECOSSOCPSolver2013], we obtain the optimal solution of the subproblem, denoted $\boldsymbol{z}^{(k)}_{\star}$. This solution serves simultaneously as the reference trajectory for the next iteration $\boldsymbol{z}_{\mathrm{ref}}^{(k+1)}=\boldsymbol{z}^{(k)}_{\star}$. The SCP forward process is terminated once the improvement in the objective falls below the prescribed threshold given by
@@ 3.2.2 Backward Process of Differentiable Sequential Convex Programming
$\mathcal{L}=\mathcal{L}\left(\boldsymbol{z}_{\mathrm{ref}}^{(1)},\dots,\boldsymbol{z}_{\mathrm{ref}}^{(K)},\boldsymbol{\theta}_{\mathrm{SCP}},\boldsymbol{\theta}_{\mathrm{TP}}\right)=\mathcal{L}\left(\boldsymbol{\theta}^{(1)},\dots,\boldsymbol{\theta}^{(K)},\boldsymbol{\theta}^{(K+1)}\right)$ (39)
where $\boldsymbol{\theta}^{(K+1)}=\left(\boldsymbol{z}_{\star}^{(K)},\boldsymbol{\theta}_{\mathrm{SCP}},\boldsymbol{\theta}_{\mathrm{TP}}\right)$. The cost function thus depends explicitly on the optimal solution of the final iteration as well as on all intermediate iterations through the sequence of reference trajectories. Since the SCP process is parameterized by $\boldsymbol{\theta}_{\mathrm{SCP}},\boldsymbol{\theta}_{\mathrm{TP}}$, the chain rule gives
$\frac{\partial\mathcal{L}}{\partial\boldsymbol{\theta}_{\mathrm{SCP}}}$ $=\sum_{k=1}^{K+1}\left(\frac{\partial\mathcal{L}}{\partial\boldsymbol{\theta}^{(k)}}\right)^{\top}\frac{\partial\boldsymbol{\theta}^{(k)}}{\partial\boldsymbol{\theta}_{\mathrm{SCP}}}$ (40)
$\frac{\partial\mathcal{L}}{\partial\boldsymbol{\theta}_{\mathrm{TP}}}$ $=\sum_{k=1}^{K+1}\left(\frac{\partial\mathcal{L}}{\partial\boldsymbol{\theta}^{(k)}}\right)^{\top}\frac{\partial\boldsymbol{\theta}^{(k)}}{\partial\boldsymbol{\theta}_{\mathrm{TP}}}$
The derivative of the cost function with respect to the final optimal solution $\boldsymbol{z}_{\star}^{(K)}$ can be computed directly from the definition of the loss, yielding $\frac{\partial\mathcal{L}}{\partial\boldsymbol{z}^{(K)}}$. For an intermediate iteration $k$, the gradient with respect to the parameter set $\boldsymbol{\theta}^{(\theta)}$ is given by
$\frac{\partial\mathcal{L}}{\partial\boldsymbol{\theta}^{(k)}}=\left(\frac{\partial\mathcal{L}}{\partial\boldsymbol{z}_{\star}^{(k)}}\right)^{\top}\frac{\partial\boldsymbol{z}_{\star}^{(k)}}{\partial\boldsymbol{\theta}^{(k)}},k\in[1,K]$ (41)
The Jacobian $\frac{\partial\boldsymbol{z}_{\star}^{(k)}}{\partial\boldsymbol{\theta}^{(k)}}$ can be obtained from the linear sensitivity equation Eq.(33). However, computing this Jacobian explicitly would require forming blockwise inverses of $\boldsymbol{H}^{(k)}$ and evaluating explicit analytic derivatives of matrices such as $\boldsymbol{A}^{(k)}$ with respect to $\boldsymbol{\theta}^{(k)}$. To avoid this computational burden, an auxiliary linear system is introduced as
@@
Here, $\boldsymbol{P}^{(k)}$ is a permutation matrix obtained using an approximate minimum-degree method, $\boldsymbol{L}^{(k)}$ is lower triangular, and $\boldsymbol{D}^{(k)}$ is a diagonal matrix. A small regularization term is added to the diagonal of $\boldsymbol{H}^{(k)}$ to ensure numerical stability. Since the coefficient matrices arising in SCP are typically large-scale and sparse, exploiting sparse symmetric matrix factorizations to solve the associated linear systems is computationally advantageous and significantly accelerates CPU-based computations. Taken together, Eq.(33) and Eq.(42) form a coupled forward-backward propagation system. Eliminating intermediate variables leads to the differential of the loss
@@
$\frac{\partial\mathcal{L}}{\partial\boldsymbol{b}_{0}}=\boldsymbol{y}_{\boldsymbol{\nu}},\,\,\frac{\partial\mathcal{L}}{\partial\boldsymbol{h}_{0+}}=\boldsymbol{y}_{\boldsymbol{\mu}},\,\,\frac{\partial\mathcal{L}}{\partial b_{i}}=y_{s,i},\,\,\frac{\partial\mathcal{L}}{\partial\boldsymbol{h}_{i}}=-\lambda_{i,0}\left(\boldsymbol{I}-\frac{\boldsymbol{g}_{i,1}\boldsymbol{g}_{i,1}^{\top}}{g_{i,0}^{2}}\right)\frac{\boldsymbol{G}_{i}\boldsymbol{y}_{\boldsymbol{z}}}{g_{i,0}}-y_{s,i}\frac{\boldsymbol{g}_{i,1}}{g_{i,0}}$
@@
Eq.(40), Eq.(45), and Eq.(46) describe the complete backpropagation pathway of the DSCP framework.
@@ 4 Simulations and Results
To demonstrate the generality and practical utility of the proposed DSCP framework, three numerical experiments are conducted across two representative trajectory planning scenarios: powered powered guidance (PDG) for reusable rockets and entry flight for hypersonic gliding vehicles (HGVs). These experiments target three different categories of parameters, including nonconvex decision variables, hyperparameters of the SCP procedure, and vehicle-design parameters. All simulations are executed in Python 3.8.20. Gradient backpropagation is implemented using PyTorch 1.12.0, and the subproblems are solved using the ECOS solver [domahidiECOSSOCPSolver2013]. As ECOS currently supports only CPU computation, all simulations are performed on the CPU within one core.
@@ 4.1 Optimization of Terminal Time in Powered Descent Guidance
This subsection applies the DSCP framework to a classical PDG problem [liFreeFinalTimeFuelOptimal2022a]. The objective is to learn and optimize the terminal time, which is a nonconvex and problem-dependent variable that influences fuel optimality.
@@ 4.1.1 Powered Descent Guidance for Reusable Rockets
$\dot{\boldsymbol{v}}=\frac{\boldsymbol{u}_{\mathrm{PDG}}}{m}+\boldsymbol{g}_{E}$
$\dot{m}=-\frac{\left\|\boldsymbol{u}_{\mathrm{PDG}}\right\|_{2}}{I_{\mathrm{sp}}g_{0}}$
The state vector is $\boldsymbol{x}_{\mathrm{PDG}}=\left[\boldsymbol{r}^{\top},\boldsymbol{v}^{\top},m\right]^{\top}$, where $\boldsymbol{r}$ and $\boldsymbol{v}$ denote the position and velocity vectors, and $m$ is the rocket mass. $\boldsymbol{u}_{\mathrm{PDG}}$ denotes the thrust of rockets. The parameter $I_{\mathrm{sp}}$ denotes the specific impulse, and $\boldsymbol{g}_{E}=[g_{0},0,0]^{\top}$ is the gravitational acceleration vector, where $g_{0}$ is the standard gravitational constant. The descent guidance is subject to the following operational constraints:
$\boldsymbol{x}_{\mathrm{PDG}}(t_{0})=\boldsymbol{x}_{\mathrm{PDG},0}$ (48)
@@
$u_{\min}\leq\left\|\boldsymbol{u}_{\mathrm{PDG}}\right\|_{2}\leq u_{\max}$
@@
$\sqrt{u_{y}^{2}+u_{z}^{2}}-\tan{\eta_{\max}}u_{x}\leq 0$
where $u_{\mathrm{min}}$ and $u_{\mathrm{max}}$ are thrust bounds, $\beta_{\max}$ is the maximum allowable glideslope angle, and $\eta_{\max}$ is the maximum thrust pointing angle. The performance index for fuel-optimal descent is
@@
Fuel-optimal powered descent is inherently a free-final-time problem. The present experiment evaluates the DSCP framework’s ability to learn and optimize such nonconvex timing parameters. The work in [liFreeFinalTimeFuelOptimal2022a] converts the free-final-time problem into multiple fixed-time problems and performs a grid search over the terminal time, followed by training a neural network to approximate the optimal relationship. In contrast, DSCP uses the gradients obtained by backpropagation through the SCP iterations, allowing direct end-to-end training of a neural network without terminal-time enumeration.
@@ 4.1.2 Verification of Differentiability
This experiment evaluates the accuracy of the gradients produced by DSCP. A virtual-control-augmented SCP method is used to solve the fixed-time fuel-optimal PDG problem. Let $K_{\mathrm{PDG}}$ be the total number of SCP iterations and $N_{\mathrm{PDG}}$ be the number of discretization intervals. The scalar cost function used for backpropagation is
@@
Simulation results are shown in Fig. 2 through Fig. 3. Figure 2(a) illustrates the relationship between terminal time and maximum remaining mass for a fixed initial condition. The results indicate that no feasible solution exists within the interval $32.00,32.05]\mathrm{s}$, and solutions in $32.05,32.10]\mathrm{s}$ exhibit abnormal behavior, where virtual control terms do not converge to zero. This failure occurs because the terminal time is too short for feasible descent. When $t_{\mathrm{f}}>32.10\mathrm{s}$, the SCP iterations converge reliably. With two-decimal precision, the optimal terminal time is $32.81\mathrm{s}$, consistent with the result reported in [liFreeFinalTimeFuelOptimal2022a].
@@ 4.1.3 The Training of Neural Network for Predicting the Terminal Time
$\boldsymbol{r}(t_{0})=[5200\pm 200,\pm 600,\pm 600]^{\top}(\mathrm{m}),\boldsymbol{v}(t_{0})=[-135\pm 15,\pm 30,\pm 30]^{\top}(\mathrm{m/s})$ (52)
@@
Figure 4 displays the training behavior, including the learning-rate profile and the average absolute value of the output gradient per epoch. As training progresses, gradient magnitudes decrease steadily, indicating improved numerical stability. Test results are presented in Figs.5 and 6. Under a representative vertical velocity-altitude profile with $r_{y}=r_{z}=100\mathrm{m},v_{y}=v_{z}=0$, the maximum relative prediction error is below $3\times 10^{-4}$. For a horizontal profile with $r_{x}=5200\mathrm{m},v_{x}=-135\mathrm{m/s},v_{y}=v_{z}=0$, the maximum relative error remains below $4\times 10^{-4}$. These results confirm that the proposed DSCP framework can be seamlessly integrated with neural networks and can effectively train them to learn variables such as the optimal terminal time in trajectory planning problems.
@@ 4.2 Trust-Region Penalty Coefficient Optimization
The trust-region penalty coefficients constitute a set of critical design parameters in construction of the convexified subproblems of SCP. This subsection considers the entry trajectory planning of HGVs and investigates the effects of the magnitudes of the trust-region penalty coefficients on the convergence rate and numerical behavior of the trajectory optimization process. In contrast to adaptive trust-region updating strategies used in previous studies, the present investigation focuses on fixed penalty values that remain constant throughout all SCP iterations.